# Fluorescence emission of the JUNO liquid scintillator

### v1

@@ Abstract
JUNO is a huge neutrino detector that will use 20 kton of organic liquid scintillator as its detection medium. The scintillator is a mixture of linear alkyl benzene (LAB), 2.5 g/L of 2,5-diphenyloxazole (PPO) and 3 mg/L of 1,4-Bis(2-methylstyryl)benzene (Bis-MSB). The main goal of JUNO is to determine the Neutrino Mass Ordering [1, 2, 3]. In order to achieve this purpose, good energy and position reconstruction is required, hence a complete understanding of the optical characteristics of the liquid scintillator is mandatory. In this paper we present the measurements on the JUNO scintillator emission spectrum, absorption length and fluorescence time distribution performed respectively with a spectrofluorimeter, a spectrophotometer and a custom made setup.
keywords:
Fluorescence , time profile , linear alkyl benzene , Cherenkov light , scintillator , neutrino physics , JUNO experiment , PPO , Bis-MSB
journal: Nuclear Instruments and Methods in Physics Research A
@@ 3.2 Measurements
The emission spectrum for the full mixture (violet solid line) peaks at 400 nm and 420 nm. The two smaller peaks at about 360 nm and 375 nm are residual of the PPO emission and in fact they dominate the LAB + 2.5 g/L PPO emission (green dashed line). It is important to notice that the emission of the JUNO mixture matches well the quantum efficiency of the 20” NNVT-PMTs that are employed in the detector (represented by the blue dots in figure 2).
@@ 5 Timing properties
The time distribution of the fluorescence light emitted by the liquid scintillator can be modelled as a superposition of different decreasing exponential distributions with characteristic time constants $\tau_{i}$ and relative weights $q_{i}$ [26]. The fastest scintillation light component from the liquid scintillator is expected to show a time constant $\tau_{1}$ at the nanosecond level, while the others range from $\sim$ few nanoseconds to hundreds of nanoseconds.
@@ 5.3 Data analysis
$F_{fit}(t)=N\sum_{i=1}^{4}\frac{q_{i}}{\tau_{i}-\tau_{r}}\left(e^{-\frac{t-t_{0}}{\tau_{i}}}-e^{-\frac{t-t_{0}}{\tau_{r}}}\right)\Theta(t-t_{0})$ (2)
where $N$ is a normalization constant, $q_{i}$ and $\tau_{i}$ are the relative weights (also called “fractions” hereafter) of the $i$-th component and its characteristic decay times, $\tau_{r}$ is the characteristic rise time (common to all the four components) and $\Theta(t-t_{0})$ is the step function that marks the start of the fluorescence time profile. The normalization of the fourth component requires a dedicated treatment due to the acquisition time window, which is limited to 1600 ns. This limitation implies a truncation in the light collection, so the normalization of the fourth exponential component $\frac{q_{4}}{\tau_{4}-\tau_{r}}$ needs to be corrected:
$\tilde{q_{4}}=\frac{1-\sum_{i=1}^{3}q_{i}}{1-e^{-\frac{t_{w}-t_{0}}{\tau_{4}}}}$ (3)
@@
To take into account the finite resolution of the setup it is necessary to convolve the fluorescence model described in equation 2 with the IRF shown in figure 7:
@@
Thanks to the linearity of the operation, the analytic convolution of the IRF with the fluorescence model in equation 2 gives:
$F_{fluo}(t)=N\sum_{i=1}^{4}\sum_{j=1}^{7}N_{i}N_{j}\left(e^{-t/\tau_{i}}-e^{-t/\tau_{r}}\right)*G_{j}(t;\mu_{j},\sigma_{j})$ (5)
@@ 5.4 Pulse Shape Discrimination performances
The fluorescence time profiles for alphas and betas shown in figure 9 are significantly different: in particular, the scintillation light induced by alpha particles has proportionally more photons emitted at longer time with respect to the one induced by beta particles. This opens the possibility to perform particle identification via the so-called Pulse Shape Discrimination (PSD) technique [20], which plays an important role for background rejection. For example, the detection of solar neutrinos via elastic scattering on electrons (which provides a beta-like signal) can greatly benefit from the rejection of alpha radioactive background. We have estimated the alpha/beta discrimination capability of the JUNO liquid scintillator by using a simple algorithm, based on the so-called “tail-to-total” ratio [^2] Note that the “tail-to-total” method is the simplest PSD technique. Better results could be achieved using more sophisticated techniques, like the Gatti or Neural Network based methods [10] $R$. $R$ is calculated for each scintillation event as the number of photons collected in the tail of the time distribution (for $t>t_{cut}$) over the total amount of photons: in average, $R$ is larger for alpha events with respect to beta events. We used a toy Monte-Carlo to simulate $10^{4}$ of alpha and beta events, following the measured profiles of figure 9, and we built the distributions of $R$. The $R$ distributions obtained with this procedure for beta and alpha events (blue dashed and red solid line respectively) are shown in figure 11. In this example, we assumed 1500 collected photon-electrons for each event corresponding to about 1 MeV of deposited energy in JUNO.
@@
$D=\frac{\mu_{\alpha}-\mu_{\beta}}{\sqrt{\sigma_{\alpha}^{2}+\sigma_{\beta}^{2}}}$ (6)

### v2

@@ Abstract
JUNO is a huge neutrino detector that will use 20 kton of organic liquid scintillator as its detection medium. The scintillator is a mixture of linear alkyl benzene (LAB), 2.5 g/L of 2,5-diphenyloxazole (PPO) and 3 mg/L of 1,4-Bis(2-methylstyryl)benzene (Bis-MSB). The main goal of JUNO is to determine the Neutrino Mass Ordering [1, 2, 3]. In order to achieve this purpose, good energy and position reconstruction is required, hence a complete understanding of the optical characteristics of the liquid scintillator is mandatory. In this paper we present the first measurements on the JUNO scintillator emission spectrum, absorption length and fluorescence time distribution performed respectively with a spectrofluorimeter, a spectrophotometer and a custom made setup of the liquid scintillator produced by the JUNO purification plants.
@@ 3.2 Measurements
The emission spectrum for the full mixture (violet solid line) peaks at 400 nm and 420 nm. The two smaller peaks at about 360 nm and 375 nm are residual of the PPO emission and in fact they dominate the LAB + 2.5 g/L PPO emission (green dashed line). It is important to notice that the emission of the JUNO mixture matches well the quantum efficiency of the 20" NNVT-PMTs that are employed in the detector (represented by the blue dots in figure 2).
@@ 5 Timing properties
The time distribution of the fluorescence light emitted by the liquid scintillator can be modelled as a superposition of different decreasing exponential distributions with characteristic time constants $\tau_{i}$ and relative weights $q_{i}$ The fastest scintillation light component from the liquid scintillator is expected to show a time constant $\tau_{1}$ at the nanosecond level, while the others range from $\sim$ few nanoseconds to hundreds of nanoseconds.
@@ 5.3 Data analysis
$F_{fit}(t)=N\sum_{i=1}^{4}\frac{q_{i}}{\tau_{i}-\tau_{r}}\left(e^{-\frac{t-t_{0}}{\tau_{i}}}-e^{-\frac{t-t_{0}}{\tau_{r}}}\right)\Theta(t-t_{0})$ (5.1)
where $N$ is a normalization constant, $q_{i}$ and $\tau_{i}$ are the relative weights (also called “fractions" hereafter) of the $i$-th component and its characteristic decay times, $\tau_{r}$ is the characteristic rise time (common to all the four components) and $\Theta(t-t_{0})$ is the step function that marks the start of the fluorescence time profile. The normalization of the fourth component requires a dedicated treatment due to the acquisition time window, which is limited to 1600 ns. This limitation implies a truncation in the light collection, so the normalization of the fourth exponential component $\frac{q_{4}}{\tau_{4}-\tau_{r}}$ needs to be corrected:
$\tilde{q_{4}}=\frac{1-\sum_{i=1}^{3}q_{i}}{1-e^{-\frac{t_{w}-t_{0}}{\tau_{4}}}}$ (5.2)
@@
To take into account the finite resolution of the setup it is necessary to convolve the fluorescence model described in equation 5.1 with the IRF shown in figure 7:
@@
Thanks to the linearity of the operation, the analytic convolution of the IRF with the fluorescence model in equation 5.1 gives:
$F_{fluo}(t)=N\sum_{i=1}^{4}\sum_{j=1}^{7}N_{i}N_{j}\left(e^{-t/\tau_{i}}-e^{-t/\tau_{r}}\right)*G_{j}(t;\mu_{j},\sigma_{j})$ (5.4)
@@ 5.4 Pulse Shape Discrimination performances
The fluorescence time profiles for alphas and betas shown in figure 9 are significantly different: in particular, the scintillation light induced by alpha particles has proportionally more photons emitted at longer time with respect to the one induced by beta particles. This opens the possibility to perform particle identification via the so-called Pulse Shape Discrimination (PSD) technique [20], which plays an important role for background rejection. For example, the detection of solar neutrinos via elastic scattering on electrons (which provides a beta-like signal) can greatly benefit from the rejection of alpha radioactive background. We have estimated the alpha/beta discrimination capability of the JUNO liquid scintillator by using a simple algorithm, based on the so-called “tail-to-total" ratio [^3] Note that the “tail-to-total” method is the simplest PSD technique. Better results could be achieved using more sophisticated techniques, like the Gatti or Neural Network based methods [10] $R$. $R$ is calculated for each scintillation event as the number of photons collected in the tail of the time distribution (for $t>t_{cut}$) over the total amount of photons: in average, $R$ is larger for alpha events with respect to beta events. We used a toy Monte-Carlo to simulate $10^{4}$ of alpha and beta events, following the measured profiles of figure 9, and we built the distributions of $R$. The $R$ distributions obtained with this procedure for beta and alpha events (blue dashed and red solid line respectively) are shown in figure 11. In this example, we assumed 1500 collected photon-electrons for each event corresponding to about 1 MeV of deposited energy in JUNO.
@@
$D=\frac{\mu_{\alpha}-\mu_{\beta}}{\sqrt{\sigma_{\alpha}^{2}+\sigma_{\beta}^{2}}}$ (5.5)